Find an equation of the line containing the points (4,9) and (-1,8)
Answer by Edwin McCravy(20086)

Find an equation of the line containing
the points (4,9) and (-1,8)

Use the slope formula:

     y2 - y1
m = —————————
     x2 - x1

where (x1, y1) = (4, 9) and (x2, y2) = (-1, 8)

     (8) - (9)     -1     1
m = ——————————— = ———— = ——— = 1/5
    (-1) - (4)     -5     5 

Now substitute in the point slope formula:

   y - y1 = m(x - x1)

   y - 9 = (1/5)(x - 4)

Multiply both sides by LCD = 5


5(y - 9) = 5(1/5)(x - 4)

 5y - 45 = 1(x - 4)

 5y - 45 = x - 4

      5y = x + 41

     -41 = x - 5y

  x - 5y = -41       

Any of those last three equations will do.

Edwin